# Hard Contacts with Soft Gradients: Refining Differentiable Simulators for Learning and Control

### v1

@@ B.3 Contacts from distance with the straight-through-trick
This effectively means that on the forward pass, we use $F_{\theta}(t,x(t))$, whereas on backward we use $\frac{\partial\tilde{F}_{\theta}}{\partial(x,\theta)}(t,x(t))$. Crucially, the derivatives are evaluated at the unmodified forward trajectory $x(t)$. Differentiation still works readily with both ODE differentiation techniques described in the previous section. In discretize-then-optimize, automatic differentiation automaticaly uses the vector-jacobian-products derived from $\tilde{F}_{\theta}$ instead of $F_{\theta}$. In optimize-then-discretize, with the straight-through-trick applied to $F$, we replace the vector-jacobian-products in the adjoint equations Eq. S6 as follows
$\tilde{a}_{x}(T)$ $=\frac{dL}{dx(T)}$ $\dot{\tilde{a}}_{x}(t)$ $=-\tilde{a}_{x}(t)^{\top}\frac{\partial\tilde{F}_{\theta}}{\partial x}(t,x(t))$ (S9)
$\tilde{a}_{\theta}(T)$ $=0$ $\dot{\tilde{a}}_{\theta}(t)$ $=-\tilde{a}_{x}(t)^{\top}\frac{\partial\tilde{F}_{\theta}}{d\theta}(t,x(t)).$ (S10)
@@
To illustrate this, we again consider a version of the billiard-toy example from Figure 8. This time, we turn gravity off and disable the table-ball contacts to isolate the ball-ball contact, which means that the ODE is linear before and after the collision. We again compute the loss and gradient over a range of initial parameters, the results are reported in Figure S3. As before, we observe that with the DiffMJX adaptive integrator, we get well-behaved gradients, but the gradient becomes zero when the two balls do not collide.
@@
Now, we activate the CFD mechanism and repeat the experiment. In this case, the gradient show oscillations due to the different stiffness settings of the forward and adjoint ODE: The solver for the forward ODE takes very large steps when no ball-ball collision happens, but the adjoint ODE incorporates the collision signal and hence is stiff. Simply unrolling the forward integration therefore makes a discretization error, resulting in the oscillations. Note, that we did not observe this phenomenon in Figure 8, as this experiment includes ball-table collisions which cause the adaptive solver of the forward ODE to take small steps even when no ball-ball collision is happening. The solution is to use optimize-then-discretize. Here, the adaptive solver for the adjoint ODE can select stepsizes according to the stiffness of the adjoint ODE. This is also confirmed by the results in Figure S3 (right).
@@
This experiments highlights an extreme corner case, and we did not observe issues in our other experiments when using discretize-then-optimize with CFD, as we typically do not have such extreme variations of the stiffness in the forward ODE. However, we believe it is beneficial to be aware of this potential caveat and how one can resolve it with the tools available in DiffMJX.
Figure S3: Optimize-then-discretize vs discretize-then-optimize in the presence of contacts-from-distance. Results are for a reduced billiard example as in Figure 8, but without table-ball contacts and gravity. Left: DiffMJX adaptive integration results in well-behaved gradients, but gradient is zero if objects do not collide. Middle: Using CFD with discretize-then-optimize results in small gradient oscillations. The adaptive integrator selects large stepsizes as no contacts are happening (forward ODE is non-stiff), whereas in the gradient computation the CFD are added (adjoint ODE is stiff). In turn, unrolling the large stepsizes cause discretization errors. Note that this phenomenon persists no matter how low the tolerance of the forward ODE solver is set, as the stepsize controller is completely unaware of the CFD mechanism that is only used in the gradient computation. Right: Using optimize-then-discretize allows the solver for the adjoint ODE to select small stepsizes when the adjoint ODE is stiff due to CFD. This fixes the gradient oscillations.
Figure S4: Jax code for minimal collision simulation. The code simulates a minimal version of a penalty-based or ideal-elastic simulator and is used in Sec. 5 to illustrate the TOI discretization errors as shown in Figure 5.
Appendix C Evaluation details and further experiments
C.1 Gradient analysis of MJX
@@
In Sec. 5, the time discretization errors arising in penalty-based simulators are first illustrated on a minimal toy example. The code to generate this toy example is shown in Figure S4. After illustrating on this toy example that time discretization errors can be mitigated by reducing the stepsize, it is shown in Figure 5 that these errors also occur in MuJoCo. Figure S5 is an extended version of Figure 5 containing all of MuJoCo’s basic shape primitives.
Figure S6 shows the error vs runtime trade-off for different integrators in MJX and DiffMJX. We observe that in the case of the cube toss, the adaptive integrators can severely reduce the runtime required to achieve gradients with low error.
@@
Figure S5: Simulation of geometric primitives thrown onto a surface using MJX or DiffMJX. Contacts at default stepsize $0.002\,\mathrm{s}$, solref=[0.005 1.0], and solimp=[0.0 0.95 0.001 0.5 2] cause MJX’s gradients of the toss distance to deviate significantly from central difference gradients, while DiffMJX maintains close agreement.
Figure S6: Error of loss and gradient vs runtime and compilation time for different integrators. The loss and gradient are computed for the cube toss shown in Fig. S5 at an initial velocity of $v_{x}=-2.0$ with contact settings solref=[0.005 1.0] and solimp=[0.0 0.95 0.001 0.5 2]. The rollout is also visualized at the top of this figure. The ground-truth gradients are computed from finite-differences using an adaptive solver with very low tolerance ($10^{-12}$) to simulate the rollout. Standard fixed-stepsize MJX integrators have red colors, fixed-stepsize integrators in Diffrax have green color, and adaptive-stepsize integrators in Diffrax have blue color (also marked by *). For adaptive stepsize control we use the Diffrax PID controller with $P=0.2,I=0.4,D=0.0$, as recommended for stiff ODEs. We observe that the pareto-front of adaptive solvers is shifted, allowing lower loss and gradient errors with less runtime. The best-performing solver is the Tsit5 solver [68], which is a 5th order explicit Runge–Kutta method with an embedded 4th order method for adaptive step sizing. This is also the default solver used in other experiments.
C.2 System identification experiments
@@
The contactnet dataset [^3] https://github.com/DAIRLab/contact-nets consists of $550$ trajectories of an acrylic cube that has been repeatedly tossed onto a wooden table. As reported in [26, 28], data has been collected at 1480 Hz and the cube’s physical parameters amount to a side length of $10\,\mathrm{cm}$, mass of $0.37\,\mathrm{kg}$, inertia of $0.0081\,\mathrm{kg\,m^{2}}$, a friction coefficient of $0.18$, and restitution of $0.125$.
Before training:
After training (starting from pre-tuned initial conditions):
Figure S7: Top: Example trajectory predictions of DiffMJX before/after training. Center: Identification of MuJoCo parameters with DiffMJX on Contactnets cube toss dataset starting from pre-tuned inital conditions. Bottom: Identification starting from random initial parameters. The cube’s geometry has a significant effect on the time of impact such that the optimization converges to a sidelength close to 100 mm to reduce the train loss. The effect of other model parameters on the system dynamics is more inter-twined. For example, a too large mass resulting in larger penetrations during impact can be compensated by increasing the constraint stiffness via solref[0] or solimp[0].
@@
The trajectories are split into segments of length five. In the supervised train loop, given a trajectory segment’s initial state, each simulator is unrolled for four steps. Subsequently, an mean square error (MSE) loss between the segment’s states and predicted states is computed. Before being fed to the loss function, each state’s quaternion is converted to a rotation matrix to avoid representation singularities negatively affecting training [73, 74]. For adaptive integration, DiffMJX is set to use “RecursiveCheckpointing” and a 5th order explicit Runge–Kutta method (“Tsit5” in Diffrax) with PID error tolerances of 1e-5. We use the Adam [31] implementation from the Optax library [63] for gradient-based optimization with parameters [b1 = 0.5, b2 = 0.9, eps = 1e-6]. MuJoCo’s parameter are set to a timestep of 0.006767 s (the data collection sampling time), iterations=4 (contact solver iterations), ls_iterations=10 (solver steps), tolerance=1e-8, impratio=1.0, solref=[0.02 1], solimp=[0.05 0.95 0.01 0.5 2], and friction solimp=[0.4 0.01 0.1]. The cube’s mass is set to the reported groundtruth value and its inertia is computed by MuJoCo’s default equal density approximation. DiffMJX uses the same MuJoCo settings with the exception of iterations=2 (adaptive integration does not require as many solver iterations). The CFD impedance function is set [0.0, 0.01, 0.01, 1.0, 4.0] that is $d_{c}=0.01$, $d_{0}=0$ with $w_{c}=1\,\mathrm{m}$.
C.2.1 Identification of additional MuJoCo parameters
In this section, we extend the experiment from Sec. 5.1 and use DiffMJX to estimate MuJoCo’s simulation parameters alongside the cube’s side length. In these experiments, the same setup is used for DiffMJX as detailed in the previous section. In comparison to the experiment in Sec. 5.1, the integration error tolerance is increased to 1e-4 to reduce computation time. Moreover, the parameters are constrained using either a softplus function or a softclip function where the softening hyper-parameter has been carefully tuned. While the models are trained on an MSE between trajectory segments, we evaluate the model on the “trajectory error” being the mean absolute error between the first forty trajectories in the dataset and DiffMJX’s trajectory prediction. In the these experiments, DiffMJX estimates the following parameters simultaneously:
  • sidelength as also estimated in Sec. 5.1,
@@ C.2.1 Identification of additional MuJoCo parameters
The optimization is started from the same pre-tuned initial conditions as used in Sec. 5.1. The prediction horizon is set to $N=10$. The training results are shown in Figure S7 (Center). Each model requires around 10 seconds of computation time per gradient step.
@@
The optimization is started with random initial parameters. The prediction horizon is set to $N=4$ such that the majority of runs require around 5 seconds for a gradient step. The training results are shown in Figure S7 (Bottom).
In both experiments, the trajectory error remains above 0.25. As noted by Parmar et al. [27], the dynamics of a hard cube exhibit a degree of chaotic behavior. Consequently, small variations in initial conditions – potentially arising from sensor noise or external disturbances such as wind – can substantially alter the cube’s trajectory after one of its corners hits the table. Nevertheless, the reduction in trajectory error from approximately 0.32 to 0.26 results in the simulated cube notably more closely matching the real-world data, as shown in Figure S7.
In both experiments, the sidelengths of most models converged closely to the groundtruth. Albeit a small offset to the groundtruth sidelength remains for the majority of the runs.
C.3 MPC experiment
@@
All MPC experiments are performed on a NVIDIA GeForce RTX 3060. In the in-hand manipulation task the runtimes for the gradient-based MPC are $2.9$h ($+2.2$h JIT) without CFD and $6.4$h ($+2.1$h JIT) with CFD. The sampling with $1024$ samples takes $35$min ($+3$min JIT). In the bionic tennis task, the runtimes for the gradient-based MPC are $9.7$h ($+1.2$h JIT) with CFD. The sampling based MPC runs for $1.4$h with $2048$ samples.

### v2

@@ C.4 Contacts from distance with the straight-through-trick
In discretize-then-optimize, automatic differentiation automaticaly uses the vector-jacobian-products derived from $\tilde{F}_{\theta}$ instead of $F_{\theta}$. In optimize-then-discretize, with the straight-through-trick applied to $F$, we replace the vector-jacobian-products in the adjoint equations Eq.˜6 as follows
$\tilde{a}_{x}(T)$ $=\frac{dL}{dx(T)}$ $\dot{\tilde{a}}_{x}(t)$ $=-\tilde{a}_{x}(t)^{\top}\frac{\partial\tilde{F}_{\theta}}{\partial x}(t,x(t))$ (8)
$\tilde{a}_{\theta}(T)$ $=0$ $\dot{\tilde{a}}_{\theta}(t)$ $=-\tilde{a}_{x}(t)^{\top}\frac{\partial\tilde{F}_{\theta}}{d\theta}(t,x(t)).$ (9)
@@
To illustrate this, we again consider a version of the billiard-toy example from Fig.˜8. This time, we turn gravity off and disable the table-ball contacts to isolate the ball-ball contact, which means that the ODE is linear before and after the collision. We again compute the loss and gradient over a range of initial parameters, the results are reported in Fig.˜19. As before, we observe that with the DiffMJX adaptive integrator, we get well-behaved gradients, but the gradient becomes zero when the two balls do not collide.
@@
Now, we activate the CFD mechanism and repeat the experiment. In this case, the gradient show oscillations due to the different stiffness settings of the forward and adjoint ODE: The solver for the forward ODE takes very large steps when no ball-ball collision happens, but the adjoint ODE incorporates the collision signal and hence is stiff. Simply unrolling the forward integration therefore makes a discretization error, resulting in the oscillations. Note, that we did not observe this phenomenon in Fig.˜8, as this experiment includes ball-table collisions which cause the adaptive solver of the forward ODE to take small steps even when no ball-ball collision is happening. The solution is to use optimize-then-discretize. Here, the adaptive solver for the adjoint ODE can select stepsizes according to the stiffness of the adjoint ODE. This is also confirmed by the results in Fig.˜19 (right).
@@
This experiments highlights a specific corner case, and we did not observe issues in our other experiments when using discretize-then-optimize with CFD, as we typically do not have such extreme variations of the stiffness in the forward ODE. However, we believe it is beneficial to be aware of this potential caveat and how one can resolve it with the tools available in DiffMJX.
Appendix D DiffMJX & CFD: The sharp bits
CFD adds additional contacts to the solver which can slow down computation.
The documentation of MuJoCo XLA [^1] https://mujoco.readthedocs.io/en/stable/mjx.html#mjx-the-sharp-bits advises to use meshes with 200 vertices or fewer due to computational limitations on the current implementation of collision detection. If MJX is used with CFD, then every contact is added to the contact solver for which $r<w_{c}$. In turn, for dense meshes and large $w_{c}$, computation time notably increases. The parameter $w_{c}$ can be made arbitrarily small such that, in its limit for $w_{c}=0$, we retain vanilla MJX with our improvements on the differentiability of the collision detector. That said, for in-hand manipulation tasks and robot locomotion tasks, the number of objects colliding at one instance is small enough to render CFD a useful tool for easing optimization with penalty-based simulators.
Adaptive integration slows down training if tolerances are set too small.
For the very large initial cube side length initial value of $140$ mm, DiffMJX saw a significant drop in computation speed forcing us to abort these runs. This is not surprising as penetrations amounting to 40% of the total cube’s width cause huge contact forces that significantly stiffen the ODE. In Appendix E.2, we reduce the integration error tolerance which allowed optimization to converge even for large initial parameters. Alternatively, one could speed up training for these too large geometry parameter initializations by first using soft impedance settings that during training becomes annealed to become increasingly stiff. The same principle applies to scenes with multiple objects: overly tight integration error tolerances can substantially slow down the simulation. As integration error scales with relative contact velocity, scenarios involving large state changes along contact normals trigger smaller step sizes. Notably, while the ball in the billiard shot example (Fig.˜8) was rolling over the table, the integration stepsize did only marginally decrease. Whereas during the collision between balls, the integrator adjusted the stepsize notably.
Finite differences can be a viable alternative for computing low-dimensional gradients.
In general, gradient computation with MJX via automatic differentiation is notably slower than running only the forward simulation. This makes zeroth-order methods such as predictive sampling favorable whenever the task can be solved with them. This is usually the case for low-dimensional systems or computing gradients of only a few parameters, e.g. compting gradients via finite differences for less then ten parameters seems to work well in practice. As sampling suffers from the curse of dimensionality, automatic differentiation for computing gradients scales to extremely high-dimensional control problems which enabled us to do gradient-based MPC on the musculoskeletal systems.
CFD should be combined with sampling to solve nonconvex optimizations.
For MPC, our CFD mechanism tends to be effective when the object needs to be pushed in a specific direction by the hand, but it performs poorly in grasping tasks. This limitation arises because, at a distance, the cumulative force arising from the hand’s CFDs conglomerate into a force that pushes objects away. However, the gradient does not encode the possibility of grasping, as this signal only emerges when the object is inside the hand. Fundamentally, this issue reflects the non-convex nature of the optimization landscape. However, we do not believe that this is a problem that should be solved using a CFD mechanism; rather, in such tasks, the best approach would be to combine CFD with sampling to overcome the non-convexity while maintaining the favorable scaling of gradients with dimensionality.
Parallelizing DiffMJX simulations.
Locomotion policies trained with reinforcement learning typically rely on domain randomization with many environments executed in parallel. DiffMJX supports such batched execution via JAX’s vmap, analogous to MJX. However, if a single policy induces large contact forces, then this may trigger very small integration steps, which results in the other simulations to wait for a slow simulation to finish. This slowdown is mitigated by increasing the integration error tolerances or by early-terminating the respective simulation once a predefined solver-step budget is exceeded. However, it is important to note that such a slowdown is only observed when many contacts are happening in a single environment. On the other hand, when few contacts are happening in each parallelized environment, there is no additional slowdown due to vmapping. This is because the selected adaptive stepsize in Diffrax is not global across the vmapped environments, it is fully vmapped and can take different values across environments. We experimentally verify this in Fig.˜20, see the caption for details.
Figure 20: We vmap the DiffMJX simulation of a ball drop across 20 different environments. In each environment, a single ball is dropped (with an initial downwards velocity and no gravity) onto a plane, from which it bounces back up. The 20 environments start with the ball at different heights, such that the contact happens at a different simulation time for each environment. We then plot the selected adaptive stepsize of the solver over the simulation time for each environment, and observe that the stepsize is indeed independent for each run, with a reduction in stepsize only occurring at the single contact in a specific environment. Therefore, when vmapping across environments, contacts in one simulation do not slow down other simulations.
Appendix E Evaluation details and further experiments
E.1 Gradient analysis of MJX
@@
In Fig.˜4, the time discretization errors arising in penalty-based simulators are first illustrated on a minimal toy example. The code to generate this toy example is shown in Fig.˜16. After illustrating on this toy example that time discretization errors can be mitigated by reducing the stepsize, it is shown in Fig.˜4 that these errors also occur in MuJoCo. Fig.˜21 is an extended version of Fig.˜4 containing all of MuJoCo’s basic shape primitives.
@@
Figure 21: Simulation of geometric primitives thrown onto a surface using MJX or DiffMJX. Contacts at default stepsize $0.002\,\mathrm{s}$, solref=[0.005 1.0], and solimp=[0.0 0.95 0.001 0.5 2] cause MJX’s gradients of the toss distance to deviate significantly from central difference gradients, while DiffMJX maintains close agreement.
E.2 System identification experiments
@@
The contactnet dataset [^2] https://github.com/DAIRLab/contact-nets consists of $550$ trajectories of an acrylic cube that has been repeatedly tossed onto a wooden table. As reported in Pfrommer et al. (2021); Acosta et al. (2022), data has been collected at 1480 Hz and the cube’s physical parameters amount to a side length of $10\,\mathrm{cm}$, mass of $0.37\,\mathrm{kg}$, inertia of $0.0081\,\mathrm{kg\,m^{2}}$, a friction coefficient of $0.18$, and restitution of $0.125$.
@@
The trajectories are split into segments of length five. In the supervised train loop, given a trajectory segment’s initial state, each simulator is unrolled for four steps. Subsequently, an mean square error (MSE) loss between the segment’s states and predicted states is computed. Before being fed to the loss function, each state’s quaternion is converted to a rotation matrix to avoid representation singularities negatively affecting training Geist et al. (2024); Brégier (2021). For adaptive integration, DiffMJX is set to use “RecursiveCheckpointing” and a 5th order explicit Runge–Kutta method (“Tsit5” in Diffrax) with PID error tolerances of 1e-5. We use the Adam Kingma & Ba (2015) implementation from the Optax library DeepMind et al. (2020) for gradient-based optimization with parameters [b1 = 0.5, b2 = 0.9, eps = 1e-6]. MuJoCo’s parameter are set to a timestep of 0.006767 s (the data collection sampling time), iterations=4 (contact solver iterations), ls_iterations=10 (solver steps), tolerance=1e-8, impratio=1.0, solref=[0.02 1], solimp=[0.05 0.95 0.01 0.5 2], and friction solimp=[0.4 0.01 0.1]. The cube’s mass is set to the reported groundtruth value and its inertia is computed by MuJoCo’s default equal density approximation. DiffMJX uses the same MuJoCo settings with the exception of iterations=2 (adaptive integration does not require as many solver iterations). The CFD impedance function is set [0.0, 0.01, 0.01, 1.0, 4.0] that is $d_{c}=0.01$, $d_{0}=0$ with $w_{c}=1\,\mathrm{m}$.
E.2.1 Identification of additional MuJoCo parameters
In this section, we extend the experiment from Section˜5.1 and use DiffMJX to estimate MuJoCo’s simulation parameters alongside the cube’s side length. In these experiments, the same setup is used for DiffMJX as detailed in the previous section. In comparison to the experiment in Section˜5.1, the integration error tolerance is increased to 1e-4 to reduce computation time. Moreover, the parameters are constrained using either a softplus function or a softclip function where the softening hyper-parameter has been carefully tuned. While the models are trained on an MSE between trajectory segments, we evaluate the model on the “trajectory error” being the mean absolute error between the first forty trajectories in the dataset and DiffMJX’s trajectory prediction. In the these experiments, DiffMJX estimates the following parameters simultaneously:
  • sidelength as also estimated in Section˜5.1,
@@ E.2.1 Identification of additional MuJoCo parameters
The optimization is started from the same pre-tuned initial conditions as used in Section˜5.1. The prediction horizon is set to $N=10$. The training results are shown in Fig.˜22 (Center). Each model requires around 10 seconds of computation time per gradient step.
@@
The optimization is started with random initial parameters. The prediction horizon is set to $N=4$ such that the majority of runs require around 5 seconds for a gradient step. The training results are shown in Fig.˜22 (Bottom).
In both experiments, the trajectory error remains above 0.25. As noted by Parmar et al. (2021), the dynamics of a hard cube exhibit a degree of chaotic behavior. Consequently, small variations in initial conditions – potentially arising from sensor noise or external disturbances such as wind – can substantially alter the cube’s trajectory after one of its corners hits the table. Nevertheless, the reduction in trajectory error from approximately 0.32 to 0.26 results in the simulated cube notably more closely matching the real-world data, as shown in Fig.˜22.
E.3 MPC experiment
@@
All MPC experiments are performed on a NVIDIA GeForce RTX 3060. In the in-hand manipulation task the runtimes for the gradient-based MPC are 2.92.9h (+2.2+2.2h JIT) without CFD and 6.46.4h (+2.1+2.1h JIT) with CFD. The sampling with $1024$ samples takes 3535min (+3+3min JIT). In the bionic tennis task, the runtimes for the gradient-based MPC are 9.79.7h (+1.2+1.2h JIT) with CFD. The sampling based MPC runs for 1.41.4h with $2048$ samples.
@@
Figure 22: Top: Identification of MuJoCo parameters with DiffMJX on Contactnets cube toss dataset starting from pre-tuned inital conditions. Bottom: Identification starting from random initial parameters. The cube’s geometry has a significant effect on the time of impact such that the optimization converges to a sidelength close to 100 mm to reduce the train loss. The effect of other model parameters on the system dynamics is more inter-twined. For example, a too large mass resulting in larger penetrations during impact can be compensated by increasing the constraint stiffness via solref[0] or solimp[0].